# On the Utility of Accounting for Human Beliefs about AI Intention in Human-AI Collaboration

### v2

@@ Abstract.
To enable effective human-AI collaboration, merely optimizing AI performance without considering human factors is insufficient. Recent research has shown that designing AI agents that take human behavior into account leads to improved performance in human-AI collaboration. However, a limitation of most existing approaches is their assumption that human behavior remains static, regardless of the AI agent’s actions. In reality, humans may adjust their actions based on their beliefs about the AI’s intentions, specifically, the subtasks they perceive the AI to be attempting to complete based on its behavior. In this paper, we address this limitation by enabling a collaborative AI agent to consider its human partner’s beliefs about its intentions, i.e., what the human partner thinks the AI agent is trying to accomplish, and to design its action plan accordingly to facilitate more effective human-AI collaboration. Specifically, we developed a model of human beliefs that captures how humans interpret and reason about their AI partner’s intentions. Using this belief model, we created an AI agent that incorporates both human behavior and human beliefs when devising its strategy for interacting with humans. Through extensive real-world human-subject experiments, we demonstrate that our belief model more accurately captures human perceptions of AI intentions. Furthermore, we show that our AI agent, designed to account for human beliefs over its intentions, significantly enhances performance in human-AI collaboration.
Key words and phrases:
Human-AI Collaboration, Reinforcement Learning, Belief Inference
footnotetext: Preprint. Under review.
1. Introduction
@@
The potential for human-AI collaboration is immense and spans various domains. In healthcare, AI systems can identify diagnoses that might be overlooked by human professionals (Cheng et al., 2016; Mobadersany et al., 2018). In industrial manufacturing, robots work alongside human workers to enhance efficiency and safety (Bauer et al., 2008; Sherwani et al., 2020). In workflow productivity, virtual assistants can generate drafts for humans to refine and finalize (Wilson and Daugherty, 2018). However, despite significant improvements in AI performance over the past decade, designing AI agents to optimize the overall performance of human-AI collaboration remains a challenge.
@@
In particular, optimizing the AI system in isolation is often shown to be insufficient for enhancing the performance of human-AI collaboration. To optimize collaborative performance, AI agents need to account for their human counterparts when deciding on their strategies. For example, in AI-assisted decision-making, researchers have demonstrated that, instead of optimizing AI independently, training AI to focus on improving in areas where humans typically struggle can significantly enhance the performance of human-AI teams (Bansal et al., 2021; Wilder et al., 2021). In human-AI collaboration, Carroll et al. (2019) show that incorporating models of human behavior into the training of AI agents leads to higher collaborative performance compared to training the agent through self-play (Silver et al., 2018). These studies highlight the importance of integrating human behavior into the design of collaborative AI. However, a key limitation of these research efforts is that they mostly assume human behavior remains static, irrespective of the actions and behavior of the AI counterpart. In practice, humans may modify their behavior in response to their beliefs about the intentions of AI agents based on their observations of AI behavior.
@@
In this work, we argue and aim to demonstrate that, in addition to incorporating human behavior, designing an AI agent that accounts for humans’ beliefs about AI intentions can significantly improve their collaborative performance. To illustrate this argument, consider a simple example task in Figure 1. In this task, both the human and AI need to collaborate to put out a fire in the center of the map. To successfully extinguish the fire, they need to use both a fire extinguisher and a hydrant, with the human and AI agent each retrieving one of them. Assuming they cannot directly communicate but can observe the direction the other is moving towards, if the AI agent believes the human will simply take the tool closest to them (the hydrant) and will not adjust their actions based on the AI’s behavior, the AI would choose to take the longer route to get the fire extinguisher. However, if the human would actually update their actions based on their inference of what the AI is trying to do from observing the AI’s actions, the more efficient strategy for the AI would be to move towards the hydrant and let the human grab the fire extinguisher. This example task demonstrates that, if humans would update their behavior based on their beliefs over AI intentions, we can design AI that account for human belief to improve overall human-AI performance.
(a) When the AI agent does not account for human beliefs about AI intentions and only considers human behavior, it would assume that the human would grab the closer hydrant. Therefore, the AI would go for the fire extinguisher, which would take 12 steps to put out the fire.
(b) When the AI agent accounts for human beliefs over AI intentions, it will try to reveal its intentions for humans so humans can take a complementary tool that would reach the overall better outcome. In this case, it takes 9 steps for the fire to be put out.
Figure 1. An example human-AI cooperation task that demonstrates the benefits for the AI to account for human beliefs about AI intentions.
In this work, we investigate and examine the benefits of incorporating human beliefs about AI intentions in the design of AI to improve human-AI collaboration. To operationalize the concept of AI intentions, our work focuses on a setting where human and AI agents need to coordinate their goals (e.g., they need to complete different subtasks) to accomplish the overall task. In this setting, the intention of the AI is defined as the goal or subtask the AI is attempting to achieve.
@@
More formally, we formulate the human-AI collaboration environment as a multi-player goal-oriented Markov decision process (MDP). We begin our investigation by developing models of human beliefs regarding AI intentions. This modeling effort extends the level-$k$ reasoning framework (Stahl II and Wilson, 1994) to account for suboptimal human behavior. Specifically, we first develop a behavioral level-0 model that assumes agents take actions without considering the behavior of other agents. We then develop the belief model by introducing a behavioral level-1 model, which assumes humans interpret the behavior of another player as if the other agent take actions according to the behavioral level-0 model. With the models of human behavior and beliefs established, we proceed to develop collaborative AI agents that incorporate different assumptions about human models.
@@
To examine our proposed approaches, we conduct extensive human-subject experiments. To develop and assess the model of human behavior (i.e., the behavioral level-00 model), we first train models of human behavior using behavioral cloning on real-world human data and evaluate their performance. For the development and assessment of the model of human beliefs about AI intentions (i.e., the behavioral level-$1$ model), we conduct experiments that present each participant with a trace of behavior by an agent and ask them to infer the agent’s goal, examining whether our model accurately predicts human inference of the agent’s intention. Moreover, based on the belief model, we explore the possibility of developing AI policies that make it easier for humans to infer the intentions of AI from its actions. Finally, utilizing the developed models of human behavior and beliefs about AI intentions, we conduct a final set of human-subject experiments in which each human participant pairs with an AI agent in a two-player coordination game. We assess the human-AI collaborative performance for different designs of AI agents. Overall, our results demonstrate the effectiveness of our developed models of human behavior and beliefs about AI intentions. Furthermore, we show that AI agents accounting for models of human behavior and beliefs about AI intentions achieve better collaborative performance when working with humans compared to AI agents that either do not consider human beliefs or disregard both human behavior and beliefs.
Contributions. In this work, we demonstrate that incorporating human beliefs about AI intentions in the design of collaborative AI can significantly improve human-AI collaborative performance. More concretely, the main contributions of our work are summarized as follows:
@@
    We develop models of human beliefs about AI intentions, inspired by the level-$k$ framework, and extend this framework to account for human suboptimality. In addition, we design AI agents capable of executing explicable action plans, making it easier for humans to infer AI intentions based on observed AI actions.
@@
    We design collaborative AI agents that incorporate both models of human behavior and human beliefs about AI intentions, aiming to optimize the performance of human-AI collaboration.
@@
    Through extensive human-subject experiments, we demonstrate that our belief models are effective and can guide the design of AI policies that are explicable to humans. Moreover, our design of collaborative AI that integrates human beliefs about AI intentions significantly improves collaborative performance when paired with real-world human participants. These results underscore both the importance of incorporating human beliefs about AI intentions in the design of collaborative AI and the effectiveness of our approach.
2. Related Work
@@
Our work contributes to the growing body of research on human-AI collaboration (Carroll et al., 2019; Jaderberg et al., 2019; Bansal et al., 2021; Wilder et al., 2021). For example, Bansal et al. (2021) demonstrate that optimizing AI in isolation may lead to suboptimal performance in human-AI collaboration, compared to training AI with consideration for human performance. Wilder et al. (2021) show that training AI systems to complement humans by performing better in areas where humans struggle results in improved collaborative performance. Carroll et al. (2019) illustrate that incorporating a human model, learned from human data, into the training of AI enhances performance when these AI systems collaborate with real humans. Our work extends this line of research by not only accounting for human behavior when designing AI but also incorporating human beliefs about AI intentions.
While these approaches require knowledge or data of humans, there have been zero-shot coordination techniques being proposed to train cooperative AI agents when human data are not available. Jaderberg et al. (2019) use population-based reinforcement learning to improve the robustness of trained AI agents. Strouse et al. (2021) develop a variety of self-play agents by varying random seeds of neural network initialization and various checkpoints, and train a best response model tailored to the agent population, discovering that humans prefer to partner with this model. Zhao et al. (2023) utilize the advances of maximum entropy reinforcement learning to encourage the diversity of agent populations and show the learned AI models could adapt to real humans. In this work, we have included these approaches as baselines and show that accounting for human beliefs lead to better performance in our setting.
From a technical perspective, our work focuses on understanding and modeling human decision-making. A significant body of literature has explored modeling human behavior. For example, inverse reinforcement learning (IRL)(Ng et al., 2000; Abbeel and Ng, 2004; Ramachandran and Amir, 2007) aims to infer reward functions in Markov decision processes (MDPs) by observing demonstrations of the optimal policy. When the demonstrator is human, these demonstrations may be noisy or exhibit behavioral biases. Several studies(Evans et al., 2016; Shah et al., 2019; Hughes et al., 2020; Zhi-Xuan et al., 2020) have addressed this by incorporating human behavioral biases into the inference process, aiming to infer both the rewards and biases simultaneously. Imitation Learning (Hussein et al., 2017) also seeks to develop models that can mimic human behavior based on demonstrations. Recently, there has been a growing amount of research that incorporates human models into computational and machine learning frameworks (Kleinberg and Oren, 2014; Tang and Ho, 2019, 2021; Masters et al., 2021a, b; Yu and Ho, 2022; Yu et al., 2023; Feng et al., 2024).
@@
From the perspective of modeling human beliefs about others’ intentions, this has been discussed in the literature on level-$k$ reasoning (Stahl II and Wilson, 1994; Gill and Prowse, 2016) in economics and in the theory of mind (Premack and Woodruff, 1978; Westby and Riedl, 2023). For instance, Yu et al. (2006) found that real human behavior aligns closely with level-$1$ and level-$2$ reasoning models in cooperative games. Albrecht and Stone (2018) survey works on autonomous agents modeling the beliefs and intentions of other agents, distinguishing between methods for modeling stationary or changing agent behaviors. These belief models include theory of mind, recursive reasoning models, plan recognition, partially observable Markov decision processes (POMDPs), and others. Baker et al. (2011) propose a Bayesian theory of mind to describe how humans model joint beliefs of state in a partially observable MDP, conducting human experiments in the works of Baker et al. (2009, 2017). Similarly, Wu et al. (2021) apply a human modeling approach in the Overcooked experiment, breaking down the task of delivering a dish into several subtasks, such as picking up an ingredient or placing an ingredient into a pot, where the player infers others’ intentions regarding these subtasks and selects their own accordingly before taking actions. Our belief inference framework differs from previous theory of mind literature by modeling and incorporating human beliefs about AI intentions specifically, using a more realistic human model as the basis for inferring these intentions.
3. Problem Formulation and Methods
In this section, we first formulate the human-AI collaboration framework as a multi-player goal-oriented Markov decision process. We then outline our methods for modeling human behavior and beliefs about AI behavior. Finally, we present our approaches for integrating human behavior and beliefs into the development of collaborative AI agents.
3.1. Decision-Making Environment
@@
We formulate the human-AI cooperative decision-making environment as a multi-agent MDP, represented by $W=\langle S,\alpha,{A_{i}|_{i\in[\alpha]}},R,P\rangle$, where $S$ is a finite set of states, $\alpha$ is a finite set of players, $A_{i}$ is the action set available to player $i$, $P:S\times A_{1}\times\dots\times A_{|\alpha|}\times S\rightarrow[0,1]$ is the transition function that determines the next state given all players’ joint actions, and $R:S\times A_{1}\times\dots\times A_{|\alpha|}\rightarrow\mathbb{R}$ is the joint reward function assigned to all players given their joint actions. Note that in this work we consider the cooperative setting. Therefore, our formulation only incorporates a single reward function $R$ for all players, though this can be easily extended.
While our formulation could address cases with multiple human and AI players, in this work, we focus on a two-player cooperative game, i.e., $\alpha={1,2}$, where one player is a human and the other is the AI. During the decision-making process, neither the human nor the AI knows the other player’s next action or future plans, and they cannot communicate directly. However, they can observe each other’s past actions, allowing them to infer the intentions of the other player and adjust their own actions accordingly.
@@
Agent intentions. In this work, we focus on a setting with a goal-oriented MDP, where there is a set of goals $G={g_{1},\dots,g_{k}}\subseteq S$. The decision-making agent only receives rewards when arriving at one of the goal states, and the goal states are terminal states, i.e., $P(g|g,a)=1\leavevmode\nobreak\ \forall g\in G,a\in A$. Both humans and AI agents can observe past actions but not the specific goals the other agents are trying to achieve. We utilize this formulation to operationalize agent intentions; that is, the goal an agent is trying to reach is defined as the agent’s intention. In the discussion of this paper, we interchangeably use both goals and intentions to refer to intentions.
3.2. Modeling Human Beliefs about AI Intention
@@
Our models are motivated by the level-$k$ framework (Stahl II and Wilson, 1994) in economics. In particular, we start by considering humans as level-00 agents who do not account for others’ behavior in the environment. Unlike previous literature, we address the realistic scenario where level-00 agents may not behave optimally, and we call this model behavioral level-00 agents. To account for human beliefs about AI intentions, [^1] In this work, we sometimes use human beliefs to refer specifically to humans’ beliefs about the intention of the AI agent. we model humans as an extension of level-$1$ agents, who assume other agents are behavioral level-00 agents [^2] Our model can iteratively extend to higher level-$k$ agents; however, we focus on the case with $k\leq 1$. and update their beliefs about AI intentions in a Bayesian manner based on the observation of others’ behavior.
@@
Modeling human behavior. We first describe our models of human behavior under the assumption that humans do not consider other players in the environment (or that they view other players as part of the environment without strategically responding). Our approach to modeling human behavior follows standard methods in the literature and primarily serves as a foundation for developing models of human beliefs. In particular, a human behavior model can be represented as $H:W\rightarrow\Pi$, mapping a given environment $w\in W$ to a policy $\pi=H(w)$. We provide two examples of human behavior models used in our work below.
@@
    Behavioral level-00 model. We also consider the case where we can learn human behavioral models from their historical behavioral traces using behavioral cloning approaches (Pomerleau, 1988; Torabi et al., 2018). Behavioral cloning is an imitation learning approach that learns a policy from human demonstrations by mapping states to actions using supervised learning methods (Bain and Sammut, 1995). We build a fully connected neural network, where the input is the state encoding and the output is the probability distribution over the action space. The model is trained using the standard gradient descent method with a cross-entropy loss.
Modeling human beliefs. We now describe how we develop models for human beliefs about AI intentions. As summarized earlier, our belief models follows the ideas of level-$k$ reasoning and assume that human decision-makers perceive other agents in the environment as behavioral level-00 agents. In our discussion, we also refer to this belief model as behavioral level-$1$ agents.
More specifically, we model the human belief updating process as a Bayesian inference process, which can be represented as in Equation (1), where $\lambda(g)$ represents the human’s prior belief distributions of AI intentions, and $Pr(s_{t},a_{t}|g)$ denotes the probability of observing $(s_{t},a_{t})$ given the intention (goal) $g$, according to the policy model $\pi(a|s,g)$. This policy model represents how humans perceive the actions of other players. If a human believes the other agent is following the standard model, then the policy is derived from the optimal policy. On the other hand, if humans believe the agent follows the behavioral cloning model, then the policy will be the output of the human model learned from data.
$B(g|(s,a)_{1:t})$ $\propto\lambda(g)Pr((s,a)_{1:t}|g)$ (1)
$=\lambda(g)\prod_{1:t}Pr(s_{i},a_{i}|g)$
$=\lambda(g)\prod_{1:t}\pi(a_{i}|s_{i},g)P(s_{i}|s_{i-1},a_{i-1})$
3.3. Designing Explicable AI Policy

### v3

@@ Abstract
To enable effective human-AI collaboration, optimizing AI performance alone is insufficient without considering human factors. Recent research has shown that designing AI agents that account for human behavior leads to improved performance in collaborative settings. However, most existing approaches assume that human behavior remains static regardless of the AI agent’s actions. In practice, humans often adjust their behavior based on their beliefs about the AI’s intentions, that is, what they believe the AI is trying to accomplish. In this paper, we develop collaborative AI agents that account for their human partner’s beliefs about the AI agent’s intentions and design their action plans accordingly. We first propose a model of human beliefs, extending the classical level-k reasoning framework, that captures how people interpret and reason about their AI partner’s intentions. Using this belief model, we develop an AI agent that incorporates both human behavior and human beliefs when planning its interactions. Through extensive human-subject experiments, we show that our belief model more accurately captures human perceptions of AI intentions. Furthermore, we demonstrate that collaborative AI agents designed to account for human beliefs about their intentions significantly improve performance in human-AI collaboration.
1 Introduction
@@
The potential for human-AI collaboration is immense and spans various domains, including healthcare (Cheng et al., 2016; Mobadersany et al., 2018), industrial manufacturing (Bauer et al., 2008; Sherwani et al., 2020), and workflow productivity (Wilson and Daugherty, 2018). However, despite significant improvements in AI performance over the past decade, designing AI agents capable of effectively collaborating with humans remains a challenge.
@@
In particular, optimizing the AI system in isolation is shown to be insufficient for enhancing the performance of human-AI collaboration. To optimize collaborative performance, AI agents need to account for their human counterparts when deciding on their strategies. For example, in AI-assisted decision-making, researchers have demonstrated that, instead of optimizing AI independently, training AI to focus on improving in areas where humans typically struggle can significantly enhance the performance of human-AI teams (Bansal et al., 2021; Wilder et al., 2021). In human-AI collaboration, Carroll et al. (2019) show that incorporating models of human behavior into the training of AI agents leads to higher collaborative performance compared to training the agent through self-play (Silver et al., 2018). These studies highlight the importance of integrating human behavior into the design of collaborative AI. However, a key limitation of these research efforts is that they mostly assume that human behavior remains static, irrespective of the actions and behavior of the AI counterpart. In practice, humans may modify their behavior in response to their beliefs about the intentions of AI agents based on their observations of AI behavior.
(a) AI ignoring human beliefs about AI intentions.
(b) AI acounting for human beliefs over AI intentions.
Figure 1: An example human-AI cooperation task.
@@
In this work, we argue and aim to demonstrate that, in addition to incorporating human behavior, designing AI agents that account for human beliefs about AI intentions can further improve collaborative performance. Consider the illustrative task in Figure 1, where a human and an AI agent must collaborate to extinguish a fire at the center. To succeed, they must use both a fire extinguisher and a hydrant, with each retrieving one. They cannot communicate directly, but they can observe each other’s movements. If the AI believes the human will simply take the closest tool (the hydrant) and will not adjust their actions based on the AI’s actions, the AI would take the longer route to the fire extinguisher. However, if the human would update their behavior based on inferred AI intentions, i.e., what the AI appears to be aiming for, the AI can adopt a more efficient strategy: move toward the hydrant and let the human retrieve the fire extinguisher. This example shows that when humans update their behavior based on their beliefs about AI intentions, an AI that accounts for those beliefs can enhance overall human-AI collaboration.
@@
We investigate and examine the benefits of incorporating human beliefs about AI intentions in the design of AI to improve human-AI collaboration. Our work focuses on a setting where human and AI agents need to coordinate their goals (e.g., they need to complete different subtasks) to accomplish the overall task. We formulate the human-AI collaboration environment as a multi-player goal-oriented Markov decision process (MDP). We begin our investigation by developing models of human beliefs about AI intentions. This modeling effort extends the level-$k$ reasoning framework  (Stahl II and Wilson, 1994) to account for suboptimal human behavior. Specifically, we first develop a behavioral level-0 model that assumes agents take actions without considering the behavior of other agents. We then develop the belief model by introducing a behavioral level-1 model, assuming humans interpret others’ behavior as if the other agent takes actions according to the behavioral level-0 model.
To develop and evaluate our model of human behavior (i.e., the behavioral level-0 model), we train behavioral cloning models on real-world human data and assess their performance. To build and test our model of human beliefs about AI intentions (i.e., the behavioral level-1 model), we conduct experiments where participants observe an agent’s behavior and infer its goal, allowing us to evaluate how well our model predicts human inferences. Using this belief model, we also explore designing AI policies that help humans more easily infer AI intentions from observed actions. Overall, our results demonstrate the effectiveness of our developed models of human behavior and beliefs.
@@
Utilizing the developed models of human behavior and beliefs, we conduct a final set of human-subject experiments in which each human participant pairs with an AI agent in a two-player coordination game. We assess human-AI collaborative performance for different designs of AI agents. We show that AI agents accounting for models of human behavior and beliefs achieve better collaborative performance when working with real-world human participants compared to AI agents that either do not consider human beliefs or disregard both human behavior and beliefs.
Contributions. The main contributions of our work are summarized as follows:
@@
    We develop models of human beliefs about AI intentions, extending the level-$k$ framework to account for human suboptimality. In addition, we design AI agents capable of executing explicable action plans, making it easier for humans to infer AI intentions based on observed AI actions.
@@
    We design collaborative AI agents that incorporate both models of human behavior and models of human beliefs about AI intentions. To the best of our knowledge, this is among the first efforts to explicitly account for human beliefs about AI intentions in the design of collaborative AI.
@@
    Through extensive human-subject experiments with roughly 1,000 participants, we demonstrate that (1) our belief models are effective and can inform the design of AI policies that are explicable to humans, and (2) our collaborative AI design significantly outperforms state-of-the-art baselines when working with real humans. These results underscore the importance of incorporating human beliefs into collaborative AI and demonstrate the effectiveness of our approach.
2 Related Work
@@
Our work contributes to the growing body of research on human-AI collaboration (Carroll et al., 2019; Jaderberg et al., 2019; Bansal et al., 2021; Wilder et al., 2021). For example, Bansal et al. (2021) show that optimizing AI in isolation can lead to suboptimal performance in human-AI collaboration. Wilder et al. (2021) find that training AI to complement humans by excelling where humans struggle improves collaborative outcomes. Carroll et al. (2019) demonstrate that incorporating a human model, learned from data, into AI training enhances collaboration with real users. Our work builds on this research by accounting not only for human behavior but also for human beliefs about AI intentions.
@@
Modeling human beliefs about others’ intentions has been extensively studied in the literature on level-k reasoning (Stahl II and Wilson, 1994; Gill and Prowse, 2016) in economics and in theory of mind (Premack and Woodruff, 1978; Westby and Riedl, 2023). For example, Yu et al. (2006) found that human behavior aligns closely with level-1 and level-2 reasoning models in cooperative games. Albrecht and Stone (2018) survey methods for autonomous agents modeling the beliefs and intentions of others, distinguishing between approaches for stationary versus dynamic behaviors. These models include theory of mind, recursive reasoning, plan recognition, and others. Baker et al. (2011) propose a Bayesian theory of mind to explain how humans form joint beliefs over states in partially observable MDPs, with supporting experiments in Baker et al. (2009, 2017). Similarly, Wu et al. (2021) apply a human modeling approach in the Overcooked experiment, decomposing the task of delivering a dish into subtasks, where players infer others’ intentions and choose actions accordingly. Our belief inference framework differs from prior theory of mind work by focusing specifically on human beliefs about AI intentions, using a more realistic human model as the basis for inferring those beliefs. A more detailed discussion of related work is provided in Appendix A.
3 Problem Formulation and Methods
3.1 Decision-Making Environment
@@
We formulate the human-AI cooperative decision-making environment as a multi-agent MDP, represented by $W=\langle S,\alpha,{A_{i}|_{i\in[\alpha]}},R,P\rangle$, where $S$ is a finite set of states, $\alpha$ is a finite set of players, $A_{i}$ is the action set available to player $i$, $P$ is the transition function that determines the next state given all players’ joint actions, and $R$ is the joint reward function assigned to all players given their joint actions. Note that in this work we consider the cooperative setting. Therefore, our formulation only incorporates a single reward function $R$ for all players, though this can be easily extended. We also focus on a two-player cooperative game, where one player is a human and the other is an AI agent. During decision making, neither the human nor the AI knows the other player’s next action or future plans, and they cannot communicate directly. However, they can observe each other’s past actions, allowing them to infer the intentions of the other player and adjust their own actions accordingly.
@@
Agent intentions. In this work, we focus on a setting with a goal-oriented MDP, where there is a set of goals $G=\{g_{1},\dots,g_{k}\}\subseteq S$. The decision-making agent only receives rewards when arriving at one of the goal states, and the goal states are terminal states, i.e., $P(g|g,a)=1~{}\forall g\in G,a\in A$. Both humans and AI agents can observe past actions but not the specific goals the other agents are trying to achieve. We use this formulation to define an agent’s intention in practical terms, that is, an agent’s intention refers to the specific goal the agent is aiming to achieve.
3.2 Modeling Human Beliefs about AI Intention
@@
Our model extends the level-$k$ framework (Stahl II and Wilson, 1994) in economics. In particular, we start by modeling human behavior by considering humans as level-00 agents who do not account for others’ behavior. Unlike prior literature, we address the realistic scenario where level-00 agents may not behave optimally, and we call this model behavioral level-00 agents. Then, to model human beliefs about others’ intentions, [^1] In this work, we sometimes use human beliefs to refer to humans’ beliefs about the intention of the AI agent. we model humans as an extension of level-$1$ agents, who assume other agents are behavioral level-00 agents [^2] Our model can iteratively extend to higher level-$k$ agents; however, we focus on the case with $k\leq 1$. and update their beliefs about AI intentions in a Bayesian manner based on the observation of others’ behavior. More details are discribed below.
@@
Modeling human behavior. We first describe our models of human behavior under the assumption that humans do not consider other players (or that they view other players as part of the environment without strategically responding). Our approach follows standard approaches in the literature and primarily serves as a foundation for developing models of human beliefs. In particular, a human behavior model can be represented as $H:W\rightarrow\Pi$, mapping a given environment $w\in W$ to a policy $\pi=H(w)$. We provide two examples of human behavior models used in our work below.
@@
    Behavioral level-00 model. We also consider the case where we can learn human behavioral models from their historical behavior using behavioral cloning (Pomerleau, 1988; Torabi et al., 2018). Behavioral cloning is an imitation learning approach that learns a policy from human demonstrations by mapping states to actions using supervised learning methods (Bain and Sammut, 1995). We build a fully connected neural network, where the input is the state encoding and the output is the probability distribution over the action space. The model is trained using the standard gradient descent method with a cross-entropy loss.
Modeling human beliefs. As summarized earlier, our belief models extend the ideas of level-$k$ reasoning and assume that human decision-makers perceive other agents in the environment as behavioral level-00 agents. We also refer to this belief model as behavioral level-$1$ agents.
More specifically, we model the human belief updating process as a Bayesian inference process, which can be represented as in Equation (1), where $\lambda(g)$ represents the human’s prior belief distributions of AI intentions, and $Pr(s_{t},a_{t}|g)$ denotes the probability of observing $(s_{t},a_{t})$ given the intention (goal) $g$, according to the policy model $\pi(a|s,g)$. This policy model represents how humans perceive the actions of other players. If a human believes the other agent is following the standard model, then the policy is derived from the optimal policy. If they believe the agent follows the behavioral cloning model, the policy will reflect the learned human model.
$B(g|(s,a)_{1:t})\propto\lambda(g)Pr((s,a)_{1:t}|g)=\lambda(g)\prod_{1:t}\pi(a_{i}|s_{i},g)P(s_{i}|s_{i-1},a_{i-1})$ (1)
3.3 Designing Explicable AI Policy